OpenStudy (crashonce):

tan 70 - tan 20 = 2cot40 @mayankdevnani help pls ive been waiting for u

OpenStudy (mayankdevnani):

you have to prove this

OpenStudy (crashonce):

ye

OpenStudy (crashonce):

any ideas lol

OpenStudy (mayankdevnani):

i proved it....

OpenStudy (mayankdevnani):

convert whole equation into cot use this formula :- \[\large \bf \tan(90-A)=cotA\] so we get, \[\large \bf \cot20-\cot70\] then use this :- \[\large \bf cotA=\frac{\cos A}{sinA}\] we get, \[\large \bf \frac{\cos20}{\sin20}-\frac{\cos70}{\sin70}\] then take LCM and simplify it, \[\large \bf \frac{\sin70\cos20-\sin20\cos70}{\sin20\sin70}\] we know that \(\large \bf sinAcosB-sinBcosA=\sin(A-B)\) Use this,we get \[\large \bf \frac{\sin(70-20)}{\sin20\sin70}\] \[\large \bf \frac{\sin50}{\sin20\sin70}\] Then,multiply and divide by 2, \[\large \bf \frac{2\sin50}{2\sin20\sin70}\] Use this, \[\large \bf 2sinAcosB=\cos(A-B)-\cos(A+B)\] we get, \[\large \bf \frac{2\sin50}{\cos50-\cos90}\] We know that cos90=0 and convert sin 50 into cos 40 by using this formula,sin(90-A)=cosA After all ,we get \[\large \bf \frac{2\cos40}{\cos50}\] convert cos50 into sin40 [ cos(90-A)=sinA ] we get, \[\large \bf \frac{2\cos40}{\sin40}\] we know that, \[\large \bf farc{\cos40}{\sin40}=\cot40\] using this,we get \[\large \bf 2\cot40\] \[\huge \cal \color{red}{Hence}~\color{Blue}{Proved} \color{green}{\smile}\]

OpenStudy (mayankdevnani):

@CrashOnce sorry...for taking too much time...

OpenStudy (mayankdevnani):

Here is my mistake: i have corrected my mistake :- we know that, \[\large \bf \frac{\cos40}{\sin40}=\cot40\]

OpenStudy (mayankdevnani):

please go through this pic

OpenStudy (anonymous):

wow amazing work @mayankdevnani I'm super impressed @-@ <3

OpenStudy (mayankdevnani):

thank you .....super impressed..lolz

OpenStudy (anonymous):

you're welcome :) you must be a math major... really this proof @-@

OpenStudy (mayankdevnani):

oh!! thanks a lot

OpenStudy (crashonce):

thanks so much!!!!!!!!!

OpenStudy (mayankdevnani):

welcome :)